Night Report 2019-08-28

Night Report

Observers: Emma Callis, Seán Brennan

Support Team: Ihanec, Gromadzki

TAT: Maria Teresa Botticella, Hanin Kuncarayakti, Janet Chen

Calibrations:

EFOSC:

Daytime -

pessto_bias.obx

pessto_calib.obx

Evening Twilight -

Skyflats UBz, ran out of twillight, will try get VRi again

Weather:

Start: Clear, seeing 0.7"

Middle: Clear, seeing 0.7" - 2"

End: Clear, seeing 1"

Comments:

Mainly using Las Campanas DIMM

Big issue with the telescope, TO saying azimuth went to standby, and when trying to initialise system there was something wrong with the encoders,

telescope wasn't reaching the initialisation point

Couldn't resolve issue tonight, will have to wait for daytime technical team

EFOSC